# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v11

@@ Abstract
This paper addresses the current lack of a unified formal framework in machine learning theory, as well as the absence of robust theoretical foundations for interpretability and ethical safety assurance. We first construct a formal information model, employing sets of well-formed formulas (WFFs) to explicitly define the ontological states and carrier mappings for the core components of machine learning. By introducing learnable and processable predicates, as well as learning and processing functions, we analyze the logical inference and constraint rules underlying causal chains in models, thereby establishing the Machine Learning Theory Meta-Framework (MLT-MF). Building upon this framework, we propose universal definitions for model interpretability and ethical safety, and rigorously prove and validate four key theorems: the equivalence between model interpretability and information existence, the constructive formulation of ethical safety assurance and two types of total variation distance (TVD) upper bounds. This work overcomes the limitations of previous fragmented approaches, providing a unified theoretical foundation from an information science perspective to systematically address the critical challenges currently facing machine learning.
Keywords: machine learning, objective information theory (OIT), formal logic systems, interpretability, ethical safety assurance, estimation of error bound
@@ 1 Introduction
Although artificial intelligence (AI), exemplified by machine learning (ML), is among the most dynamic areas in information science and technology, several critical shortcomings persist in the current research landscape. First, there is a lack of a unified logical framework that spans the entire lifecycle of machine learning. Most existing ML theory focuses on isolated components: optimization theory emphasizes training dynamics, generalization analysis targets test performance, while initialization strategies and deployment are often considered independently. This fragmentation precludes the coherent characterization of knowledge evolution from prior injection, data-driven updates, to task-level generalization (Vapnik, 1999; Alquier, 2024; Schölkopf et al., 2021; Garcez and Lamb, 2023; Marra et al., 2024).
Second, the unpredictability of model performance has led to rapidly escalating research and development costs. As model sizes continue to grow, the demands for data, computation, energy, time, and human resources have multiplied. In the absence of explicit causal semantics, structural analysis, and effective metrics, it is often challenging to analyze or predict attainable performance in each stage of the pipeline based on error upper bounds. Incomplete information coverage, suboptimal model selection, inaccurate measurement, and quality deficiencies frequently result in failed development efforts and significant resource losses (Achiam et al., 2023; Touvron et al., 2023; Chowdhery et al., 2023; Gunasekar et al., 2023).
Third, there are profound concerns regarding ethical safety assurance. As AI capabilities advance, so too do apprehensions about their ethical implications. However, there is still no consensus—either mathematical or theoretical—on the scientific definition, substantive content, forms of expression, or criteria for ethical safety in AI. Most research remains focused on ethical principles and regulatory mechanisms, while the proliferation of competing guidelines leaves practitioners uncertain about which standards to follow (Russell et al., 2015; Burr and Leslie, 2023; Tilala et al., 2024; Hawkins et al., 2021; Pant et al., 2024).
Fourth, the theoretical foundations for interpretability remain weak. The lack of a unified semantic framework for causal inference and constraint logic means that traditional ML methods treat the transformation from data to model as a black box. Although interpretability is widely recognized as a critical issue and diverse approaches have been proposed to address it, there is as yet no rigorous definition or universally applicable evaluative criteria (Lipton, 2018; Glanois et al., 2024; Zhang et al., 2024; Rudin, 2019; Miller, 2019).
As a central discipline bridging data, knowledge, and systems, information science has begun to offer new perspectives for the theoretical integration of machine learning (Shwartz-Ziv and Tishby, 2017; Xu et al., 2022). However, information science theory remains fragmented across information theory, control theory, signal processing, statistical learning, and computational theory, each developing largely in isolation and lacking a unified semantic and logical system. This fragmentation is a primary reason for the absence of a unified theoretical framework in contemporary AI. In response, Objective Information Theory (OIT) applies an axiomatic approach to construct a sextuple model of information, representing intrinsic meaning via ontological states, objective expression via carrier states, and revealing the dual-subject structure of information through the enabling mapping from the former to the latter. OIT thus provides a general and universal foundational model and metrics for understanding information in the objective world, and especially within computing and information systems (Xu et al., 2014; Xu, 2024). For example, (Xu et al., 2025) demonstrates that applying OIT’s general information metrics to pre-evaluate training data can significantly enhance model training efficiency.
This paper further defines the enabling mapping, clarifying its dual mathematical and physical-causal attributes. Through a recursively defined higher-order logical system, we formalize the expression of states via the interpretation of sets of well-formed formulas (WFFs), elucidating the mathematical essence of state representation in information definitions. This fills the final theoretical gap in OIT and offers a more powerful toolkit for applying OIT to diverse problems in information science. The introduction of noisy information and information chains further establishes the theoretical groundwork for analyzing key challenges across all stages of machine learning.
On this basis, we comprehensively apply OIT’s theories and methodologies—centered on the learnable and processable predicates, as well as learning and processing functions—to five canonical stages in machine learning: model construction, training information, learning and training, input of problems, information processing outputs and online learning. We establish a unified semantic formalization of state representation, and develop causal-chain logical inference and constraint rules covering the full ML lifecycle, culminating in the Meta-Framework for Machine Learning Theory (MLT-MF).
Applying MLT-MF, we formulate universal definitions for model interpretability and ethical safety, and prove theorems relating interpretability to information existence, as well as the theoretical underpinnings for ethical safety assurance. For model error, we establish total variation distance (TVD) upper bounds for two strategies—conservative prediction in regions not covered by training with noiseless input, and arbitrary prediction with noisy input—providing new guidelines for the estimation and optimization of machine learning performance. Finally, we illustrate and validate the practical utility of MLT-MF and the associated theorems with a simple feedforward neural network model, demonstrating that this research offers a novel direction and methodology for both theoretical analysis and engineering practice in machine learning, grounded in a unified formal logical framework.
@@ 2 Related Work
2.1 Conceptual Representations of Machine Learning
Generally, machine learning refers to methods that learn general rules from limited observational data and use these rules to predict unknown information. Numerous works have provided more detailed analyses of this concept. Representative views include:
In the context of statistical learning (Vapnik, 1999), the learning model is decomposed into three components: a set of independent random vectors $x$ that follow a certain but unknown probability distribution $P(x)$; a set of output vectors $y$ corresponding to input $x$, also following a certain but unknown conditional probability distribution $P\left(y\middle|x\right)$; and a collection of functions $f(x,\alpha),\alpha\in\Lambda$, that the learning machine can implement. Machine learning is thus defined as selecting, based on a training dataset composed of $l$ independent observations $\left(x_{1},y_{1}\right),\cdots,(x_{l},y_{l})$ drawn from the same distribution $P(x,y)=P(x)P\left(y\middle|x\right)$, a function $f(x,\alpha)$(with $\alpha\in\Lambda$) that minimizes the risk function $R(\alpha)=\int{L(y,f(x,\alpha))d}P(x,y)$, where $L(y,f(x,\alpha))$ is the loss function of $f(x,\alpha)$.
Another perspective highlights the central role of model parameters (Alquier, 2024), describing machine learning as the process of learning from an object-label set $(X,Y)$ governed by a probability distribution $P$, to produce a family of predictors $\left\{f_{\theta},\theta\in\Theta\right\}$ parameterized by $\theta\in\Theta$. The loss function $L(f_{\theta}(X),Y)$ quantifies the performance of each predictor, and the expected risk is defined as $R(\theta)=\mathbb{E}_{(X,Y)\sim P}\ [L(f_{\theta}(X),Y)]$. The ultimate goal is to find $\theta^{*}$ such that $R\left(\theta^{*}\right)=\inf_{\theta\in\Theta}R(\theta)$.
From a more advanced perspective, the focus shifts to causal relationships beyond data distribution features (Schölkopf et al., 2021). This view treats perceptual data acquired through measurement as entangled observations of unknown causal system states. It seeks to learn concrete representations that at least partially reveal these hidden causal structures by leveraging large-scale training, data augmentation, self-supervised learning, and robust fine-tuning using data from multiple simulated environments. This includes discovering causal models implicit in environmental data and performing causal inference to plan and act based on these models.
Some researchers emphasize the integration of neural and symbolic artificial intelligence (Garcez and Lamb, 2023), arguing that knowledge should be rooted in efficient data learning via message passing in neural networks and embedded in vector representations as efficient computational models. Symbols, in turn, should be acquired through querying and extracting knowledge from trained networks, providing a rich descriptive language at appropriate levels of abstraction—allowing finite means to achieve infinite ends—while enabling compositional and discrete reasoning at the symbolic level, thereby permitting extrapolation beyond data distributions. The combination of learning and reasoning should offer an important alternative for tackling compositional reasoning problems, reducing the number of effective combinations through learning and thus generating simpler symbolic descriptions within a neuro-symbolic cycle.
Another viewpoint proposes that structure learning through parameter learning is the most prominent paradigm among the statistical relational learning approaches within neuro-symbolic AI (Marra et al., 2024). The learning task involves estimating probabilities for a fixed set of clauses, with the advantage of removing combinatorial search from the learning process.
The above works collectively reflect, from different perspectives, the academic community’s comprehensive understanding of both past achievements and future directions in machine learning. However, none have established a formalized expression for the entire lifecycle implementation process of machine learning under a unified semantic framework. As a result, it remains consistently difficult to apply structured systems analysis methods to guide the implementation and optimization of models throughout research and design.
2.2 Model Interpretability
The interpretability of machine learning has attracted significant attention in recent years. As noted in (Lipton, 2018), few authors in the academic literature are able to precisely articulate what interpretability actually means, or to clearly explain how their proposed solutions achieve it. Despite the absence of a unified definition, an increasing number of publications continue to introduce algorithms purported to possess interpretability. This situation supports two possible conclusions: first, that there is a universally accepted understanding of interpretability that simply remains unwritten; or second, that the term itself is inherently ambiguous, leading to quasi-scientific claims regarding the interpretability of various models. In reality, the objective of interpretability may simply be to extract more valuable information from models.
(Glanois et al., 2024) argues that, in the context of reinforcement learning, what is considered understandable and what constitutes a good explanation may differ entirely depending on whether the end-user, system designer, or policymaker is being considered. Based on a definition of comprehensibility encompassing simulatability, decomposability, and algorithmic transparency, the authors emphasize that interpretability in reinforcement learning relies on three pillars: interpretable inputs, interpretable models, and interpretable decisions. They further analyze the specific mechanisms for achieving interpretability in each of these aspects.
In (Zhang et al., 2024), the impact of different training mechanisms on two key factors—model smoothness and gradient similarity—is investigated from the perspective of agents, with the aim of deepening our understanding of the dynamics of model transferability and its relationship to various training methods. Building on adversarial training, two working hypotheses are proposed and validated: first, that the trade-off between model smoothness and gradient similarity largely determines the adversarial transferability of agent models; and second, that shifts in data distribution can undermine gradient similarity, thereby resulting in generally poorer substitutes. This approach—explaining changes in model performance by dissecting and adjusting key factors in the training process—should be regarded as an important direction in the study of machine learning interpretability.
It is evident, therefore, that although interpretability is a critical issue in ML, the lack of a unified semantic and formal framework for machine learning has precluded the establishment of rigorous definitions and universally applicable criteria for interpretability. Approaches to enhancing interpretability remain diverse, and there is still no universally accepted methodology informed by a common theoretical foundation.
@@ 2.3 Ethical Safety Assurance
Concerns regarding the ethical safety of AI have a long history. As noted in (Russell et al., 2015), the proliferation of intelligent and autonomous systems has given rise to significant legal and moral questions, the answers to which impact both producers and consumers of AI technologies. These issues span law, public policy, professional ethics, and philosophical morality, necessitating the expertise of computer scientists, legal scholars, political scientists, and ethicists. While mathematical tools such as formal logic, probability theory, and decision theory have provided important foundations for reasoning and decision-making, numerous challenges remain unresolved.
Building on traditional concepts of assurance, (Burr and Leslie, 2023) offers a definition of ethical assurance for AI: it is a process that employs structured argumentation to provide auditable guarantees—specifically, that a set of normative claims regarding a system’s ethical properties are justified in light of the available evidence. Accordingly, the general structure of moral assurance is characterized as an iterative process of reflection, action, and adjustment, repeated throughout the design, development, and deployment lifecycle.
In the context of AI/ML-supported healthcare, ethical safety assurance faces more concrete challenges. (Tilala et al., 2024) provides a comprehensive exploration of the multifaceted ethical considerations in deploying AI in healthcare. The authors advocate for multiple strategies, including the implementation of rigorous data security protocols, prioritization of data minimization, increasing diversity in training datasets, use of data augmentation to address imbalance, application of algorithmic fairness techniques to detect and mitigate model bias, regular auditing of learning algorithms, enhancing the interpretability of algorithmic outputs, regulatory oversight based on standardized performance metrics and stringent testing protocols, and the augmentation—rather than replacement—of clinicians’ capabilities as means to achieve ethical safety.
(Hawkins et al., 2021) further investigates safety assurance guidelines for machine learning in specific domains such as autonomous systems, establishing a causal chain from ML safety to software safety and ultimately to system safety. The authors propose that ML safety requirements should be derived from the system-level safety requirements, and that these should be evaluated in light of the allocated safety needs, system and software architecture, and operational environment. This framework then serves as the basis for developing safety assurance specifications across all stages of the ML lifecycle, including data preparation, model training, model validation, and model deployment.
(Pant et al., 2024) conducts an extensive study of the technology practitioners who have the greatest influence on AI ethical safety, proposing a classification framework that encompasses practitioner awareness, perceptions, needs, challenges, and approaches. The study highlights that, in the absence of a universally adopted set of ethical principles for AI, incorporating morality into AI is fundamentally challenged by the lack of clear moral knowledge and the ambiguity of ethical principles.
It is thus evident that, despite widespread concerns over the ethical safety of machine learning, there is as yet no mathematical or theoretical consensus regarding its scientific definition, substantive content, forms of manifestation, or criteria for assessment. Instead, ethical safety remains an open-ended concept, explored through various technical, engineering, and moral perspectives. Most research either focuses on specific ethical principles and consequences for AI, or on methods for enhancing ethical conduct. The proliferation of often conflicting AI ethical guidelines has, paradoxically, made it difficult for practitioners to determine which standards to follow. This underscores the necessity of a universally accepted definition of ethical safety, as such a definition can bridge the gap between ongoing academic discourse and industry practice, thereby providing a concrete foundation for strengthening ethical safety assurance in AI.
@@ 2.4 Model Error Estimation
Model performance is fundamentally reflected in the errors it may incur on specific tasks. (Vapnik, 1999) provides key theorems on the consistency of empirical risk minimization (ERM); it characterizes the conditions for both two-sided and one-sided uniform convergence from empirical risk to true risk using VC entropy, VC dimension, and growth functions. Based on VC dimension, it further establishes distribution-independent risk upper bounds and sufficient conditions for fast convergence under both bounded and unbounded loss in finite-sample settings. By controlling the trade-off between VC dimension capacity and empirical risk through nested hypothesis families, it proves the strong consistency and convergence rates of structural risk minimization (SRM). These theoretical foundations have led to practical implementations such as support vector machines (SVM), margin maximization, and kernel methods, offering explanations for their generalization mechanisms.
(Rudin, 2019) replaces global complexity with local empirical Rademacher averages, measuring complexity on subsets with small empirical error. It provides finite-sample error upper bounds via the fixed point of sub-root functions, thus achieving optimal or near-optimal convergence rates inversely proportional to the square root of the number of training samples. Under controlled variance conditions, it presents a unified framework for error estimation, including both distribution-dependent bounds determined by local Rademacher complexity and data-dependent bounds constructed from empirical local Rademacher averages. These results are extended to typical classification and regression tasks, providing computable upper bounds and iterative algorithms for quickly approximating the fixed point.
From an information-theoretic perspective, (Miller, 2019) models learning algorithms as random channels mapping input datasets $S$ to output hypothesis spaces $W$. For sub-Gaussian (including bounded) losses and uncountable hypothesis spaces, the generalization error upper bound is characterized via the mutual information $I(S;W)$ between input and output. Based on this, high-probability sample complexity and expected generalization error upper bounds are derived, and the concept of information stability $(\varepsilon,\mu)$-stability is introduced, unifying algorithmic consistency and overfitting control through information leakage measures. The paper further analyzes two classes of algorithms—Gibbs learning and noisy ERM—where $I(S;W)$ is explicitly controlled via mutual information regularization, and discusses methods for decomposing and additively controlling mutual information in complex learning algorithms through adaptive combinations.
(Shwartz-Ziv and Tishby, 2017) provides a foundational theoretical framework for meta-reinforcement learning, focusing on in-distribution cross-task generalization and training convergence. It introduces a task distribution T and task-level MDPs, defines generalization error as the performance gap between training and unseen tasks under a meta-policy, and establishes generalization upper bounds determined by the number of training tasks and task complexity, using empirical process theory and concentration inequalities. The analysis further discusses the impact of algorithmic capacity and learning rates on error bounds. Within the non-convex SGD framework, sufficient conditions and rates for finite-step convergence of meta-learning loss are presented, along with guidelines for long-term behavior and hyperparameter selection based on linear and sub-linear types. The theoretical results are validated and the observability of convergence rates is demonstrated through controllable task generation and baseline meta-algorithms.
(Xu et al., 2022) addresses the separation and convergence complexity of continuous and discrete errors in deterministic sampling of generative models. From both theoretical and numerical perspectives, it investigates the convergence properties of deterministic samplers based on probability flow ordinary differential equations (ODEs). Assuming access to score function estimates with $L_{2}$ accuracy, it provides upper bounds on TVD between target and generated distributions in continuous time. For the practical implementation using Runge-Kutta integrators, discrete-time error upper bounds are established. Numerical experiments on 128-dimensional problems are conducted to validate these theoretical error bounds.
Collectively, these works encompass both classical statistical learning theory and recent advances in reinforcement learning and generative models, reflecting a broad recognition of the importance of model error upper bounds in machine learning. However, these studies do not employ explicit causal semantic modeling, nor do they address the full lifecycle challenges of multi-stage information transmission, online learning, or multi-agent integration in ML. Additionally, they rarely consider the geometric scale and entropy of distribution shifts and uncovered domains, making it difficult to offer actionable engineering or data-sampling guidance for test domain coverage gaps and noisy input scenarios.
3 Preliminary Theory
Undoubtedly, as a key component of artificial intelligence, modern machine learning is a product of the rapid development and high integration of information technologies such as computers, the Internet, big data, and cloud computing. Machine learning systems are, without question, an important branch of information systems. Starting from the basic concept of information can provide a unified theoretical foundation for research in machine learning and artificial intelligence.
3.1 Definition of Information
(Xu et al., 2014), based on four fundamental postulates, proposes that information is an enabling mapping from the state of an ontology to the state of a carrier, but does not further elaborate on the notions of “state” or “enabling mapping.” Appendix A establishes a higher-order formal system $\mathcal{L}$ and, drawing on general common sense, formulates basic postulates regarding states. It is then proven that any set of states can be represented as the interpretation of a set of WFFs in $\mathcal{L}$ (Qiu and Xu, 2025), indicating that both ontological states and carrier states correspond to interpretations of sets of WFFs over specific domains of objects and times. Here, we proceed to further clarify the concept of enabling mapping.

### v12

@@ Abstract
This paper addresses the current lack of a unified formal framework in machine learning theory, as well as the absence of robust theoretical foundations for interpretability and ethical safety assurance. We first construct a formal information model that employs sets of well-formed formulas (WFFs) to explicitly define the ontological states and the carrier mappings of core ML components. By introducing learnable and processable predicates, as well as learning and processing functions, we analyze the logical inference and constraint rules underlying causal chains in models, thereby establishing the Machine Learning Theory Meta-Framework (MLT-MF). Building upon this framework, we propose universal definitions for model interpretability and ethical safety, and rigorously prove four key theorems: (i) the equivalence between model interpretability and information existence; (ii) a constructive formulation of ethical safety assurance; and (iii–iv) two upper bounds on total variation distance (TVD).This work overcomes the limitations of previous fragmented approaches, providing a unified theoretical foundation from an information science perspective to systematically address the critical challenges currently facing machine learning.
keywords:
machine learning, objective information theory (OIT), formal logic systems, interpretability, ethical safety assurance, error bound estimation
journal: Information Sciences
[…]
organization=Koguan School of Law, China Institute for Smart Justice, School of Computer Science, Shanghai Jiao Tong University, city=Shanghai, postcode=200030, country=China
@@ 1 Introduction
Machine learning (ML) has progressed rapidly, yet several foundational gaps persist. The field still lacks a unified logical framework that spans the full lifecycle from prior specification and data-driven learning to deployment and online adaptation. Existing theory remains fragmented across training dynamics, initialization, test-time generalization, and deployment, which prevents a coherent account of knowledge evolution across stages [1, 2, 3, 4, 5]. Model performance is also hard to predict, which inflates development costs in data, computation, energy, and human effort. In the absence of explicit causal semantics, structural analysis, and stage-wise error metrics, attainable performance levels and error upper bounds are difficult to estimate, and information coverage gaps or failure modes due to model selection, measurement, and data quality are hard to anticipate [6, 7, 8, 9]. Ethical safety lacks a settled mathematical definition and verification protocol. Current research is dominated by principles and heterogeneous guidelines, leaving uncertainty about standards and assurance processes [10, 11, 12, 13, 14]. Interpretability similarly lacks a rigorous, domain-agnostic foundation. Without a unified semantic framework for causal inference and constraint logic, the data-to-model pipeline is treated as a black box, and general, testable definitions and criteria remain elusive despite numerous approaches [15, 16, 17, 18, 19].
Information science offers a route to unification [20, 21], but its subfields—information theory, control, signal processing, statistical learning, and computation—evolved largely in isolation and lack a common semantics. Objective Information Theory (OIT) addresses this by axiomatizing information as an enabling mapping from ontological states to carrier states. Intrinsic meaning is captured by ontological states and objective expression by carrier states, and the mapping has both mathematical and physical–causal attributes [22, 23]. Prior work shows that OIT’s general metrics can improve data evaluation and training efficiency in machine learning [24]. We complete a remaining gap by formalizing state as an interpretation of well-formed formulas (WFFs) in a higher-order logical system, and by clarifying that enabling mappings are surjective mathematical mappings that are also physically realizable causal relations. We further introduce noisy information and information chains, which support analysis of multi-stage pipelines and their causal dependencies.
Building on this foundation, we develop the Machine Learning Theory Meta-Framework (MLT-MF), a unified information–logical semantics for the ML lifecycle. Our main contributions are:
• Unified lifecycle semantics. We formalize five canonical stages—model construction, training information, learning and training, problem input, and processing/output with online learning—via state objects and enabling mappings that compose into information chains.
• Logical definitions with verifiable criteria. We give domain-general, testable definitions of interpretability and ethical safety within the same semantics. We prove a necessary and sufficient condition linking interpretability to information existence along the chain and construct a maximal safe-output (SafeMax) mechanism validated via a hypergraph independent-set formulation.
• Error control with observable quantities. We derive explicit total variation distance (TVD) upper bounds in two regimes: conservative prediction in regions not covered by training under noise-free input, and arbitrary prediction under noisy input. The bounds are driven by measurable masses—overlap, uncovered, and noise (loss/addition)—and align with engineering indicators such as coverage and Kullback–Leibler (KL) divergence [1, 25, 26, 27, 28, 29, 30, 31, 32].
• Practical instantiation. On a simple feedforward network, the framework yields multi-modal interpretability through mathematical, visual, and partition-based explanations, guarantees ethical safety via SafeMax, and performance estimates via TVD bounds, illustrating deployable assurance across stages.
The remainder of the paper is organized as follows. Section 2 reviews related work on learning theory, interpretability, ethical safety, and error estimation, and situates OIT as a unifying lens. Section 3 presents theoretical basis: definition of information, noisy information, information chains, and the finite-automaton view. Section 4 introduces MLT-MF, defining stages, predicates, and functions, and specifying lifecycle constraints and inference rules with a minimal pipeline instantiation. Section 5 provides formal definitions and theorems for interpretability and ethical safety, the SafeMax mechanism, and the TVD bounds. Section 6 reports the neural-network case study and observable metrics. Section 7 discusses limitations and future directions.
@@ 2 Related Work
2.1 Conceptualizations of Machine Learning
Classical statistical learning theory formalizes learning as risk minimization over function classes under unknown data-generating distributions, with foundational tools such as VC theory, growth functions, and distribution-independent bounds [1]. Parameterized predictors, loss-based risk, and empirical risk minimization provide the standard optimization lens [2]. Recent perspectives emphasize causal structure beyond observed distributions, seeking representations that expose latent causal factors to support robust transfer and planning across environments [3]. Neuro-symbolic approaches pursue integration of distributed representations with symbolic abstraction and compositional reasoning, aiming to combine efficient learning with discrete inference [4, 5]. These lines of work clarify individual components of the learning pipeline, yet they do not supply a unified semantic and causal framework that spans model construction, training, input, inference, and deployment.
2.2 Interpretability
Despite extensive interest, interpretability lacks a precise and generally accepted definition and evaluation protocol. Prior analyses highlight ambiguity in terminology and goals, and suggest that the practical objective is to extract useful information from models in ways that are audience- and task-sensitive [15]. In reinforcement learning, interpretability depends on interpretable inputs, models, and decisions, and on notions such as simulatability, decomposability, and algorithmic transparency [16]. Mechanistic studies relate training mechanisms to properties such as smoothness and gradient similarity to explain transferability and adversarial behavior, illustrating how dissecting key factors in training can yield explanatory leverage [17]. Overall, current methods are diverse and often model- or context-specific, and they lack a unifying logical semantics that ties inputs, internal states, and outputs to causal explanations across the lifecycle [15, 16, 17, 18, 19].
@@ 2.3 Ethical Safety Assurance
As AI systems permeate high-stakes domains, ethical safety has drawn sustained attention in law, policy, and engineering. Existing work emphasizes assurance processes grounded in structured arguments that justify normative claims with evidence across the system lifecycle [10, 11]. In healthcare, recommended practices include data minimization, diversity and augmentation, algorithmic fairness, auditing, interpretability, and standards-based oversight, typically framed as a toolbox of technical and procedural controls [12]. Safety guidance for autonomous and software-intensive systems links ML safety to software and system safety via requirement derivation and lifecycle-wide assurance activities [13]. Practitioner studies report gaps in clear moral knowledge and the ambiguity of principles, which complicate operationalization and adoption [14]. What remains missing is a mathematically grounded definition of ethical safety with constructible guarantees and verifiable criteria that integrate with ML pipelines.
@@ 2.4 Model Error Estimation
Foundational results establish consistency and finite-sample bounds for empirical and structural risk minimization through VC dimension and related complexity measures, including conditions for one- and two-sided uniform convergence and margin-based generalization [1]. Local Rademacher complexity sharpens bounds by tailoring complexity to low-error regions and yields optimal or near-optimal rates via fixed-point analyses under variance control [25]. Information-theoretic approaches bound generalization via mutual information between data and hypotheses, introduce information stability, and analyze Gibbs and noisy ERM with mutual-information regularization [26]. In meta-reinforcement learning, generalization across tasks and convergence under nonconvex optimization are characterized using empirical process theory and concentration tools, with guidance for capacity and learning-rate choices [27]. For generative modeling, recent analyses provide TVD bounds for deterministic samplers via ODE perspectives and establish discrete-time error under Runge–Kutta schemes [28]. These advances quantify error in specific settings but do not explicitly model causal semantics across multi-stage information flows, nor do they integrate noise structure, coverage geometry, and uncovered-region strategies into a single computable bound that can guide data auditing and prediction strategy design.
2.5 Toward a Unified Information-Theoretic and Logical Perspective
Information science offers a unifying lens, yet its subfields have developed largely in silos [20, 21]. OIT axiomatizes information as an enabling mapping from ontological to carrier states with dual mathematical and physical–causal attributes, and provides general metrics that have shown value for data evaluation in ML [22, 23, 24]. By formalizing states as interpretations of WFFs in a higher-order logical system and introducing notions of noisy information and information chains, OIT supports analysis of multi-stage pipelines and causal dependence structures [22, 23, 33, 34]. The present work builds on this foundation to offer a single formal semantics for lifecycle reasoning about interpretability, ethical safety, and computable TVD-based error control.
3 Theoretical Basis
Modern ML systems are information systems and, in the mathematical sense, finite automata. We ground our meta-framework in a formal account of information and state that supports causal analysis across stages.
3.1 Information as Enabling Mapping
[23] based on four fundamental postulates, proposes that information is an enabling mapping from the state of an ontology to the state of a carrier, but does not further elaborate on the notion of “enabling mapping”. Here, we proceed to further clarify the concept.